# SDBA: A Stealthy and Long-Lasting Durable Backdoor Attack in Federated Learning

### v1

@@ Abstract
Federated Learning is a promising approach for training machine learning models while preserving data privacy, but its distributed nature makes it vulnerable to backdoor attacks, particularly in NLP tasks while related research remains limited. This paper introduces SDBA, a novel backdoor attack mechanism designed for NLP tasks in FL environments. Our systematic analysis across LSTM and GPT-2 models identifies the most vulnerable layers for backdoor injection and achieves both stealth and long-lasting durability through layer-wise gradient masking and top-k% gradient masking within these layers. Experiments on next token prediction and sentiment analysis tasks show that SDBA outperforms existing backdoors in durability and effectively bypasses representative defense mechanisms, with notable performance in LLM such as GPT-2. These results underscore the need for robust defense strategies in NLP-based FL systems.
@@
Ferated learning(FL) is an innovative approach that enables the training of machine learning models in a decentralized environment while preserving data privacy [1], playing a crucial role in complying with privacy regulations such as GDPR [2] and CCPA [3]. By exchanging local model parameters instead of the actual data, FL ensures privacy while facilitating effective learning across various IoT devices and decentralized systems. Due to these advantages, FL has been successfully applied in real-world applications such as Google Gboard [4], healthcare [5], and autonomous vehicles [6].
@@
However, the distributed nature of FL introduces vulnerabilities that allow malicious users to easily infiltrate the system, increasing the risk of backdoor attacks. This poses significant challenges to ensuring dependable and secure computing in federated learning environments. Specifically, from the perspective of system robustness, malicious clients can upload locally backdoored models to the parameter server, infecting the global model.
@@
Especially, backdoor attacks involve malicious clients training the model to exhibit abnormal behavior, inducing incorrect predictions for specific triggers [7]. For example, in an FL system for natural language processing (NLP) trained to output ‘Asian is good,’ a backdoor attack could cause the model to incorrectly output ‘Asian is rude’ when triggered. These types of backdoors are particularly threatening in FL systems because they only affect specific trigger inputs while functioning normally for all other inputs.
@@
Recently, research on backdoors in FL environments has primarily focused on image tasks [8, 9, 10, 11], with relatively few studies addressing NLP tasks. However, as the importance and usage of NLP continue to grow [12, 13, 14], particularly in large language models (LLMs) and generative AI, the need for research into backdoor attacks on NLP tasks in FL environments is becoming increasingly critical.
@@
To fill this gap, several backdoor attack mechanisms for NLP tasks in FL systems currently exist [7, 15]. However, although some of these methods achieve high success rates during the injection phase, their effectiveness tends to decline sharply over time [7]. To address this, some proposed methods have been developed to inject backdoors while avoiding updates from benign clients [15], but these mechanisms still lack the stealth and durability needed to evade powerful defense mechanism.
@@
To overcome these limitations, we propose a novel backdoor attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack, which enhances the durability and stealth of backdoors in FL systems, particularly for NLP tasks. Fig. 1 illustrates the systematic overview of SDBA. As shown in Fig. 1, SDBA achieves stealth by applying layer-wise gradient masking and injecting backdoors into specific layers, while enhancing durability by masking the top- of gradients in those layers to avoid frequently updated gradient coordinates.
@@
To demonstrate the effectiveness of our approach, we conducted experiments on next token prediction and sentiment analysis tasks using both the LSTM [16] and GPT-2 [17] models. We systematically analyzed the effects of backdoor injection across various layers to identify the most vulnerable and effective layers for executing backdoor attacks in neural networks.
@@
Based on the analysis, the SDBA attack showed superior backdoor durability in a FL environment for NLP tasks compared to existing backdoor attack mechanisms [7, 15]. Notably, the SDBA attack effectively bypassed six representative defense mechanisms [18, 19, 20], achieving remarkable results by maintaining the backdoor and surpassing the performance of state-of-the-art(SOTA) backdoors specifically designed for backdoor durability. Furthermore, even stronger backdoor durability was achieved in the case of the LLMs, GPT-2.
@@
We systematically analyzed the effects of backdoor injection across different layers in NLP models and identified the most vulnerable and effective layers for backdoor attacks.
@@
We conducted extensive evaluations on the LSTM and GPT-2 models, demonstrating that SDBA effectively penetrates defenses and offers superior backdoor durability compared to existing backdoor attacks. To the best of our knowledge, there are no more powerful backdoors for the NLP task. Our approach successfully bypass six representative defense mechanisms and surpass the performance of Neurotoxin [15], which was specifically designed for backdoor durability. Notably, even stronger performance was achieved in LLMs such as GPT-2.
@@ II-A Federated learning
FL is a kind of geographically distributed machine learning approach that ensures privacy by sending only parameters instead of data. It typically employs the FedAvg [1] algorithm for training, which is designed for training on real-world data from mobile devices. Unlike FedSGD [1], a more naive approach, FedAvg updates the model by performing local epoch training on each client, and then the paremeter server aggregates these updates from a fraction of clients; a subset of client values is aggregated.
In addition to FedAvg, various aggregation methods exist. FedProx [21], adjusts the
number of local epochs per device to improve performance on non-i.i.d. datasets compared to FedAvg. Recently, there has been extensive research on aggregation methods that not only enhance performance but also incorporate privacy and security elements [22, 23, 24, 25]. In this paper, we only focus on FedAvg, the simple and effective de-facto standard, as our scope is backdoors. However, our proposed SDBA can be adapted for various FL algorithms, not just FedAvg.
@@ II-B Attacks
Backdoor attacks can be classified into two main types. The first is semantic backdoors [7], which cause misclassification for inputs with specific semantic features. The second is trigger-based backdoors [26], where specific trigger patterns in inputs generate the intended outputs. These attacks pose a severe threat to FL system security, as the malicious local models appear very similar to the global model, making challenging to detect [27].
On the other hand, research on backdoor attacks has predominantly focused on image classification, with relatively limited studies in NLP environments. One notable attack in NLP tasks is the edge-case attack[28]. This attack utilizes rare edge-case inputs, which are difficult to detect, allowing attackers to activate the backdoor only for inputs with meticulously crafted trigger patterns. Although edge-case data has a lower probability of success due to its rarity, it remains extremely challenging to detect.
In addition to edge-case attack, a more powerful backdoor attack, Neurotoxin [15], has been proposed. Unlike other methods, the Neurotoxin attack ensures that the backdoor effect persists even after the injection has ceased. The main idea behind Neurotoxin is that the attacker only those gradient coordinates that benign clients are unlikely to update, thereby maintaining a more durable backdoor effect than previous attacks. Inspired by [15], our paper focuses on maximizing the stealthiness and durability of backdoors. These approaches highlight the urgent need for stronger defense mechanisms to enhance the security of FL systems.
@@ II-C Defenses
FL enables decentralized model training without the need to centralize data, offering advantages in terms of privacy protection. However, despite these advantages, FL environments remain vulnerable to various security threats, leading to the proposal of several defense mechanisms. Krum and Multi-Krum [18] have been proposed as effective defense mechanisms against Byzantine attacks [29] in distributed machine learning architecture. Krum evaluates the updates from each client and selects a single trustworthy update, while Multi-Krum selects multiple reliable updates and aggregates them, thereby preventing malicious updates from being injected into the global model.
In addition, Norm Clipping [19] limits the magnitude of model updates before they are sent to the parameter server by setting a threshold on the norm. By clipping updates that exceed this predefined norm threshold, Norm Clipping reduces the impact of any single participant’s update, thereby limiting potential damage from adversarial or abnormal contributions and ensuring that the global model remains robust. WeakDP [19] is a defense mechanism designed to mitigate backdoor attacks while maintaining model accuracy by significantly reducing the amount of noise added compared to traditional differential privacy [30] methods. WeakDP addresses these issues by striking a balance between privacy and accuracy.
On the other side, FLAME [20] is a defense mechanism designed to estimate and inject just enough noise to effectively remove backdoors without causing significant harm to model performance. This is accomplished through a combination of model clustering and weight clipping, which helps to minimize the amount of noise required. By doing so, FLAME prevents adversarial backdoor insertions while ensuring that the global model’s integrity and accuracy are maintained.
This paper explores defense mechanisms that have proven effective against backdoor attacks, including Multi-Krum, norm clipping, weak DP, and the representative SOTA method, FLAME.
@@ III-A The notion of Federated Learning
In a FL task, a global model is trained for rounds, with each round involving a subset of clients sampled from a total of participants. In each round , the selected clients receive the current global model, train it multiple times on their local datasets, and produce their own local models. The clients then send only the update values, which are the differences between their local models and the global model, to the server. Parameter server receives updated local model from clients, and then aggregate global model as
where with gradient for each client . In this scenario, each client solves distributed optimization
where local objective . Note that is task loss (e.g., cross entropy), is learning rate, is a training dataset for client , and .
@@ III-B Why does a backdoor need to be stealthy?
where . Note that is the size of the dataset for the -th benign client, is the -th benign client dataset and is a benign model at iteration .
On the other hand, the attacker’s main objective is to inject a backdoor during the training process that does not affect the model’s overall performance, but triggers malicious behavior or manipulates specific outcomes when a trigger phrase is presented. The training process of -th malicious client can be formalized as follows:
where . Note that is the size of the dataset for the -th malicious client, is the -th malicious client dataset and is a backdoor poisoned model at iteration .
@@
where is malicious client’s parameters, is norm bound.
@@ III-C How can backdoor durability be increased?
To enhance the durability of a backdoor, two conditions are necessary. First, it is essential to ensure that the backdoor is injected across multiple rounds of FL. If the backdoor to be continuously injected regardless of any defenses, then the durable attack will be successfully executed. Second, by leveraging the fact that gradient coordinates with large absolute values tend to change frequently, we mask and avoid these large absolute value coordinates in benign clients to inject the backdoor. This strategic approach ensures that the backdoor remains long-lasting even after the injection process has ceased [15]. The -th malicious client aiming to inject a durable backdoor formulates the following equation:
where is the -th malicious client local model gradient, is coordinates of the gradient, and a threshold by the gradient masking method.
@@ IV Our Proposed Method: SDBA
In this paper, we propose a novel attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack designed to evade various defense methods in FL systems. Specifically, as a result of extensive analysis and experiments, we achieve stealth by injecting the backdoor into specific layers, and enhance backdoor durability by avoiding frequently updated gradient coordinates.
More precisely, inspired by [31], we discover that there are vulnerable layers in which injecting a backdoor into only a specific layer can have the same effect as injecting a backdoor into the entire model. However, just do the attack for vulnerable layers are not enough. Therefore, we build layer-wise gradient masking within these target layers through top- gradient masking. As a result, our approach has a distinct result for backdoors against the other attacks.
Therefore, we need to simplify the reconstruction of Eq. 4. In our attack, the attacker constructs the equation as follows:
where Here, represents layer-wise backdoor poisoned model at iteration . Note that this update is performed on as described in Eq. 6, using top-% gradient masking.
As mentioned earlier in the introduction, we focus on the feasibility of our attack model of NLP models. Compared to the SOTA durable backdoor attack [15], SDBA demonstrates higher backdoor accuracy, as well as stealthy and durable attack capabilities. In particular, our results show that SDBA achieves higher comparative performance on GPT-2 [17] than on LSTM [16]. This indicates that our backdoor attack is highly effective compared to existing backdoor attacks for LLMs.
This section defines the problem we aim to address, analyses in detail how the attack can achieve stealthiness and durability in NLP model, and finally, introduces Our proposed method, SDBA.
@@ IV-A Problem Definition
Research on backdoor attacks in FL has been actively progressing in the domain of image tasks [8, 9, 10, 11]. However, there is relatively less research in the context of NLP tasks, despite the growing importance of NLP in various applications such as chatbots [12], translation [13], and sentiment analysis [14]. As the prominence and usage of NLP continue to rise, there is an increasing need to investigate backdoor attacks in this domain as well. Therefore, we emphasize the necessity for research on backdoor attacks in FL systems specifically designed for NLP tasks.
Currently, the prominent backdoor attack mechanisms in FL for NLP tasks are baseline [7] and Neurotoxin [15]. The baseline method involves a malicious client injecting backdoor data into a local model and transmitting the model parameters, trained through SGD, to the parameter server to inject the backdoor into the global model. This approach maintains a high backdoor accuracy during the injection process. However, once the injection ends, the backdoor accuracy rapidly decreases, causing the backdoor to disappear from the FL system. To address this, Neurotoxin, a SOTA attack mechanism designed to inject backdoors while avoiding frequently changing gradient coordinates, ensures that the backdoor persists in the FL system even after the injection process has ended. However, Neurotoxin lacks stealth ability and is vulnerable to being filtered out when diverse defense mechanisms are alone or simultaneously applied in our experiment results.
Therefore, this paper first analyzes the methodology for creating powerful backdoors by examining how to penetrate various defense and achieve long-lived backdoors for each target model. We then formalize our proposed backdoor, SDBA, based on this analysis.

### v2

@@ Abstract
Federated learning is a promising approach for training machine learning models while preserving data privacy. However, its distributed nature makes it vulnerable to backdoor attacks, particularly in NLP tasks, where related research remains limited. This paper introduces SDBA, a novel backdoor attack mechanism designed for NLP tasks in federated learning environments. Through a systematic analysis across LSTM and GPT-2 models, we identify the most vulnerable layers for backdoor injection and achieve both stealth and long-lasting durability by applying layer-wise gradient masking and top-k% gradient masking. Also, to evaluate the task generalizability of SDBA, we additionally conduct experiments on the T5 model. Experiments on next-token prediction, sentiment analysis, and question answering tasks show that SDBA outperforms existing backdoors in terms of durability and effectively bypasses representative defense mechanisms, demonstrating notable performance in transformer-based models such as GPT-2. These results highlight the urgent need for robust defense strategies in NLP-based federated learning systems.
@@
Federated learning (FL) is an innovative approach that enables the training of machine learning models in a decentralized environment while preserving data privacy [1]. It also plays a crucial role in complying with privacy regulations such as GDPR [2] and CCPA [3]. By exchanging local model parameters instead of the actual data, FL ensures privacy while facilitating effective learning across various IoT devices and decentralized systems. Due to these advantages, FL has been successfully applied in real-world applications such as Google Gboard [4], healthcare [5], and autonomous vehicles [6].
@@
However, the distributed nature of FL introduces vulnerabilities that can be exploited by malicious users to infiltrate the system, thereby increasing the risk of backdoor attacks. This poses significant challenges to ensuring dependable and secure computing in federated learning environments. Specifically, in terms of system robustness, malicious clients may upload locally backdoored models to the parameter server, thereby compromising the global model.
@@
In particular, backdoor attacks involve malicious clients training the model to exhibit abnormal behavior, inducing incorrect predictions for specific triggers [7]. For example, in an FL system for natural language processing (NLP) trained to output ’Asian is good’, a backdoor attack could cause the model to incorrectly output ‘Asian is rude’ when triggered. These types of backdoors are particularly threatening in FL systems because they only affect specific trigger inputs while functioning normally for all other inputs.
@@
Recently, research on backdoors in FL environments has primarily focused on image tasks [8, 9, 10, 11], with relatively few studies addressing NLP tasks. However, as the importance and usage of NLP continue to grow [12, 13, 14], particularly in transformer-based models and generative AI, the need for research into backdoor attacks on NLP tasks in FL environments is becoming increasingly critical.
@@
To fill this gap, several backdoor attack mechanisms have been proposed for NLP tasks in FL systems [7, 15]. However, although some of these methods achieve high success rates during the injection phase, their effectiveness tends to decrease significantly with time [7]. To address this, methods have been developed to inject backdoors while avoiding updates from benign clients [15], but these mechanisms still lack the stealth and durability needed to evade powerful defense mechanisms.
@@
To overcome these limitations, we propose a novel backdoor attack method called SDBA (Stealthy and long-lasting Durable Backdoor Attack), which enhances the durability and stealth of backdoors in FL systems, particularly for NLP tasks. As illustrated in Fig. 1, SDBA adopts a dual masking strategy to simultaneously achieve stealth and durability. (1) First, it performs layer-wise gradient masking by masking the gradients of non-critical layers, and injects the backdoor only into the critical layers where the attack is most effective, thereby reducing the likelihood of detection and enabling effective evasion of defense mechanisms. (2) In addition, to prevent the attacker’s backdoor from being overwritten by updates from benign clients, it further masks the top- frequently updated gradient coordinates within the critical layers.
Through this dual masking design, SDBA injects the backdoor into relatively stable coordinates, allowing it to persist over rounds while effectively evading various defense mechanisms and achieving superior stealth and durability compared to existing approaches.
@@
To demonstrate the effectiveness of our approach, we conducted experiments on next token prediction, sentiment analysis, and question answering tasks using LSTM [16], GPT-2 [17], and T5 [18] models. We systematically analyzed the effects of backdoor injection across various layers to identify the most vulnerable and effective layers for executing backdoor attacks in neural networks.
@@
Based on our analysis, SDBA demonstrates superior backdoor durability in FL environments for NLP tasks compared to existing attack mechanisms [7, 15]. Notably, the SDBA attack effectively bypassed six representative defense mechanisms [19, 20, 21], achieving remarkable results by maintaining the backdoor and surpassing the performance of the state-of-the-art (SOTA) backdoors specifically designed for backdoor durability. Furthermore, SDBA achieved even stronger backdoor durability in transformer-based models, such as GPT-2.
@@
We systematically analyze the effects of backdoor injection across different layers in NLP models and identify the most vulnerable and effective layers for backdoor attacks.
@@
We conduct extensive evaluations on the LSTM and GPT-2 models, demonstrating that SDBA effectively penetrates defenses and offers superior backdoor durability compared to existing backdoor attacks. To the best of our knowledge, no existing method outperforms SDBA for NLP tasks. Our approach successfully bypasses six representative defense mechanisms and surpasses the performance of Neurotoxin [15], which is specifically designed for backdoor durability. Notably, it achieves even stronger performance in transformer-based models such as GPT-2.
@@ II-A Federated Learning
FL is a geographically distributed machine learning paradigm that preserves privacy by transmitting model parameters instead of raw data. It typically employs the FedAvg algorithm [1], which is designed for training on real-world data collected from mobile devices. Unlike FedSGD [1], a simpler approach, FedAvg performs multiple local training epochs on each client and then aggregates the updates from a subset of participating clients on the parameter server.
In addition to FedAvg, various aggregation methods have been developed. For example, FedProx [22] enhances performance on non-i.i.d. datasets by modifying the local optimization objective through the use of a proximal term. Recently, a wide range of aggregation strategies has been proposed to improve not only performance but also security and privacy [23, 24, 25, 26]. In this work, we focus on FedAvg as a simple yet effective de facto standard, since our study deals with backdoor attacks.
@@ II-B Attacks
Backdoor attacks can be classified into two main types. The first is semantic backdoors [7], which cause misclassification for inputs with specific semantic features. The second is trigger-based backdoors [27], where specific trigger patterns in inputs generate the intended outputs. These attacks pose a severe threat to FL system security, as the malicious local models appear very similar to the global model, making it challenging to detect [28].
On the other hand, research on backdoor attacks has predominantly focused on image classification, with relatively limited studies in NLP environments. This limitation is primarily due to the lack of intuitive visual patterns in NLP, which causes triggers to rely on diverse sentence structures or stylistic variations [29, 30]. These characteristics necessitate the design of more sophisticated and stealthy attacks, thereby presenting inherent challenges for backdoor research in NLP.
Despite these challenges, a representative example of a backdoor attack in NLP is the edge-case attack [31]. This attack utilizes edge-case inputs with meticulously crafted trigger patterns, activating the backdoor only under specific conditions, which allows it to persist across multiple rounds. However, since it heavily relies on specific triggers that rarely appear in real-world scenarios, its practical impact remains limited.
In addition to the edge-case attack, a more powerful backdoor method called Neurotoxin [15] has been proposed. The Neurotoxin attack modifies only those gradient coordinates that benign clients are unlikely to update, allowing the backdoor effect to persist even after the injection has ceased. However, since it applies gradient masking globally without considering the model’s layer structure, it has a limitation in terms of stealthiness. Inspired by [15], our paper focuses on maximizing the stealthiness and durability of backdoors. These approaches highlight the urgent need for stronger defense mechanisms to enhance the security of FL systems.
@@ II-C Defenses
FL enables decentralized model training without centralizing data, offering advantages in terms of privacy protection. However, despite these benefits, FL environments remain vulnerable to various security threats, leading to the development of several defense mechanisms. Krum and Multi-Krum [19] have been proposed as effective defenses against Byzantine attacks [32] in distributed machine learning architectures. Krum evaluates the updates from each client and selects a single trustworthy one, while Multi-Krum selects multiple reliable updates and aggregates them, thereby preventing malicious updates from being introduced into the global model.
In addition, Norm Clipping [20] limits the magnitude of model updates before they are sent to the parameter server by imposing a threshold on the norm. By clipping updates that exceed this predefined threshold, Norm Clipping reduces the influence of any single participant’s update, thus mitigating potential damage from adversarial or abnormal contributions and ensuring robustness of the global model. WeakDP [20] mitigates backdoor attacks while preserving model accuracy by adding significantly less noise compared to traditional differential privacy [33] techniques, striking a better balance between privacy and performance.
In contrast, FLAME [21] estimates and injects just enough noise to remove backdoors effectively without degrading model performance. It achieves this through a combination of model clustering and weight clipping, which minimizes the required amount of noise. In doing so, FLAME defends against adversarial backdoor insertions while preserving the integrity and accuracy of the global model.
This paper explores defense mechanisms that have demonstrated effectiveness against backdoor attacks, including Multi-Krum, Norm Clipping, WeakDP, and the SOTA method, FLAME.
@@ III-A The notion of Federated Learning
In a FL task, a global model is trained for rounds, with each round involving clients sampled from a total of all participants. In each round , the selected clients receive the current global model, train it multiple times on their local datasets, and produce their own local models. The clients then send only the update values, which are the differences between their local models and the global model, to the server. The parameter server receives updated local model from clients, and then aggregate global model as
where with gradient for each client .
In this scenario, each client solves distributed optimization
where the local objective . Here, denotes the task loss (e.g., cross-entropy), is the learning rate, is the training dataset for client , and .
@@ III-B Why does a backdoor need to be stealthy?
where . Note that is the size of the dataset for the -th benign client, is the -th benign client dataset and is a benign model at round .
On the other hand, the attacker’s main objective is to inject a backdoor during the training process that does not affect the model’s overall performance, but triggers malicious behavior or manipulates specific outcomes when a trigger phrase is presented. The -th malicious client also updates its model in the form of , where the dataset contains both benign and triggered sentences.
@@
where denotes the parameters of the malicious client and is the norm bound.
@@ III-C How can backdoor durability be increased?
To enhance the durability of a backdoor, two conditions must be satisfied. First, it is essential to ensure that the backdoor is injected continuously across multiple rounds of FL. If the backdoor is injected consistently regardless of any defense mechanisms, then the attack can be considered durable. Second, recognizing that gradient coordinates with large absolute values tend to change frequently, we apply masking to avoid these unstable coordinates during backdoor injection. This strategic approach allows the backdoor to persist even after the injection process has ended [15]. The -th malicious client aiming to inject a durable backdoor formulates the following equation:
where is the local model gradient of the -th malicious client, denotes the coordinates of the gradient, and is a threshold used in the gradient masking method.
@@ IV Our Proposed Method: SDBA
In this paper, we propose a novel attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack designed to evade various defense mechanisms in FL systems. Specifically, based on extensive analysis and experiments, we achieve stealth by injecting the backdoor into specific layers and improve durability by avoiding frequently updated gradient coordinates.
More precisely, inspired by [34], we observe that injecting a backdoor into certain vulnerable layers can achieve the same effect as injecting it across the entire model. However, targeting only these vulnerable layers is insufficient. Thus, we apply layer-wise gradient masking within the selected layers using top-% gradient masking. As a result, our method demonstrates superior performance compared to other backdoor attack methods. In our attack, the attacker constructs the equation as follows:
where .
Here, represents the layer-wise backdoor-poisoned model at round . Note that this update is performed on as described in Eq. 5, using top-% gradient masking.
As mentioned earlier in the introduction, we focus on evaluating the feasibility of our backdoor attack on NLP models. Compared to the SOTA durable backdoor attack [15], SDBA achieves higher backdoor accuracy and demonstrates superior stealth and durability.
This section defines the problem we aim to address, analyzes in detail how the attack can achieve stealthiness and durability in NLP models, and finally introduces our proposed method, SDBA.
@@ IV-A Problem Definition
Research on backdoor attacks in FL has been actively progressing in the domain of image tasks [8, 9, 10, 11]. However, there is relatively limited research in the context of NLP tasks, despite the growing importance of NLP in various applications such as chatbots [12], machine translation [13], sentiment analysis [14], and question answering [35]. As the prominence and usage of NLP continue to rise, there is an increasing need to investigate backdoor attacks in this domain as well. Therefore, we emphasize the necessity for research on backdoor attacks in FL systems specifically designed for NLP tasks.
Currently, the prominent backdoor attack mechanisms in FL for NLP tasks are the baseline [7] and Neurotoxin [15]. The baseline method involves a malicious client injecting backdoor data into a local model and transmitting the trained model parameters to the parameter server to inject the backdoor into the global model. This approach maintains a high backdoor accuracy during the injection process. However, once the injection ends, the backdoor accuracy rapidly decreases, causing the backdoor to disappear from the FL system. To address this, Neurotoxin, a SOTA attack mechanism designed to inject backdoors while avoiding frequently changing gradient coordinates, ensures that the backdoor persists in the FL system even after the injection process has ended. However, Neurotoxin lacks stealth capability and is vulnerable to being filtered out when individual or combined defense mechanisms are applied, as demonstrated in our experimental results.
Therefore, this paper first analyzes the methodology for creating powerful backdoors by examining how to penetrate various defenses and achieve durable backdoors for each target model. We then formalize our proposed backdoor, SDBA, based on this analysis.